# A Large Recurrent Action Model: xLSTM enables Fast Inference for Robotics Tasks

### v1

@@ C.1 Training Tasks
C.2 Hold-out Tasks
In Figure 15, we show the zero-shot evaluation performance on the hold-out tasks 15. We want to highlight, that the performance declines for all methods and model sizes compared to performance on training tasks. This is because, hold-out tasks exhibit severe shifts in state-spaces, action-spaces and reward functions.
@@ C.2 Hold-out Tasks
C.3 Fine-Tuning
In Figure 16, we present the fine-tuning evaluation performance on the held-out tasks. We compare an xLSTM trained from scratched against an xLSTM initialized with the pre-trained weights. We use full fine-tuning for our experiments, but highlight that parameter-efficient approaches are attractive alternatives [Hu et al.,, 2022; Liu et al.,, 2022; Paischer et al.,, 2024]. We do observe consistent improvement of the pre-trained models over the models trained from scratch. This indicates that fine-tuning performance is not negatively impacted by the recurrent backbone. While we train on a substantial number of environments, the total amount of data used is, however, still only a fraction of that employed in training other large-scale models, such as LLMs. Consequently, we do not observe comparable few-shot generalization. We anticipate that few-shot generalization capabilities will emerge as we increase both data volume and model size.
@@ C.3 Fine-Tuning
C.4 In-context Learning
We assess the ICL abilities of modern recurrent architectures on the Dark-Room environment considered in prior works on in-context RL [Laskin et al.,, 2022; Lee et al.,, 2023; Kirsch et al.,, 2023; Sinii et al.,, 2023; Huang et al.,, 2024; Schmied et al., 2024b, ]. In Dark-Room, the agent is located in a dark room. The task is to navigate to an invisible goal location in that dark room. The state is partially observable, as the agent only observes its own x-y position on the grid ($|\mathcal{S}|=2$). The action space consists of 5 discrete actions: move up, move down, move left, move right, stay ($|\mathcal{A}|=5$). Upon reaching the goal location, the agent receives a reward of +1 for every step in the episode it resides on the goal location. Consequently, the agent first has to explore the room to find the goal. Once the goal location is found (as indicated by the positive reward), the agent can exploit this knowledge. Given a multi-episodic context, the agent should be able to exploit information contains in the previous trials (e.g., exploiting one path vs. avoiding another).
In our experiments, the Dark-Room is a $10\times 10$ grid and episodes last for 100 steps, starting in the top left corner of the grid. We adopt the same experiment setup as Schmied et al., 2024b and leverage their datasets. We train 16M parameter agents on datasets from 80 randomly selected goal locations in the grid. The datasets contain 100K transitions per task and are obtained by training task-specific PPO [Schulman et al.,, 2018] agents. Then, we evaluate the in-context abilities of our agents on 20 hold-out goal locations. During evaluation, the agent is given 40 episodes to interact with the environment, which we refer to as ICL-trials. Furthermore, we adopt the AD [Laskin et al.,, 2022] framework for training our agents with a multi-episodic context. We use the same sequence representation as used in our main experiments, consisting of states, returns-to-go (target return set to 80 during evaluation), and rewards. Note that this differs from the sequence representation used by Laskin et al., [2022]. We set the context length for all agents to the equivalent of two episodes, which amounts to 200 timesteps in total.
In Figure 17, we report the ICL performance over the 40 ICL trials on (a) 80 training locations and (b) 20 hold-out locations for the 4 different backbones considered in this work. We observe that the recurrent backbones attain considerably higher scores than the Transformer backbone. Furthermore, we find that xLSTM [7:1] attains the highest overall scores, which we attribute to the state-tracking abilities [Merrill et al.,, 2024] of sLSTM blocks. We aim to explore the ICL abilities of modern recurrent backbones on more complex benchmarks such as XLand-minigrid [Nikulin et al.,, 2023, 2024] in future work.
@@ C.4 In-context Learning
C.5 Inference Time Comparisons
We empirically examine the difference in inference speed between of our models. Similar to De et al., [2024], we report both latency and throughput. For real-time applications, latency is the more important dimension, and therefore we focus our analysis on latency.
C.5.1 Latency
In Figures 18 and 19, we report the latencies for DT and xLSTM with the same number of layer blocks as DT, and twice the number of layers blocks as DT, respectively. We conduct our comparison for two different batch sizes and across varying sequence lengths.
[…]
@@ C.5.1 Latency
C.5.2 Throughput
In Figures 20 and 21, we similarly report the attained throughput for DT and xLSTM with the same number of layer blocks as DT, and twice the number of layers blocks as DT, respectively. We conduct our comparison for two fixed context lengths and varying batch sizes.
[…]
@@ C.5.2 Throughput
C.5.3 xLSTM Kernel Comparisons
@@
Appendix D Ablations
D.1 Removing action condition
We found that removing actions from the context results in better performance across backbones. In Figure 23, we report the learning curves over 200K updates for DT with varying context lengths, both with and without actions in the context. While context lengths beyond 1 hurt performance on meta-world when training with actions, the reverse is true when training without actions. This is in contrast to recent works, which did not benefit from longer contexts [Octo Model Team et al.,, 2024]. However, while removing actions improves performance on the robotics domains, it does not affect performance on discrete control. For robotics, we observed that the models become overly confident (high action logits), which is problematic if poor initial actions are produced. We assume this is because in robotics actions change smoothly and by observing previous actions the agent learns shortcuts. Thus, removing actions from the input prevents the agent from using shortcuts.
@@ D.1 Removing action condition
D.2 Effect of Dropout in DT
@@
DTs use by default a Dropout [Srivastava et al.,, 2014] rate of 0.1. However, during our experiments, we found that Dropout has detrimental effects on the evaluation performance, particularly on continuous control domains like Composuite. In Figure 24, we show the validation perplexities and evaluation performance for a DT trained with and without Dropout. Consequently, we remove Dropout from our DT variant.
@@
D.3 Effect of reducing number of layers in xLSTM
In prior works, xLSTM and Mamba use twice the number of layers blocks as the Transformer baseline, while maintaining the same hidden dimension [Gu and Dao,, 2023; Beck et al.,, 2024]. For our inference-time comparisons, we therefore reduce the number of layer blocks in xLSTM by half. To ensure a fair comparison, we consequently adjust the hidden size of xLSTM to match the number of parameters of the Transformer baseline. In this section, we investigate the effect of these modifications of the xLSTM architecture on the model performance.
@@ D.3 Effect of reducing number of layers in xLSTM
Appendix E Embedding Space Analysis

### v2

@@ D.1 Training Tasks
D.2 Hold-out Tasks
In Figure 14, we show the zero-shot evaluation performance on the hold-out tasks 14. We want to highlight, that the performance declines for all methods and model sizes compared to performance on training tasks. This is because, hold-out tasks exhibit severe shifts in state-spaces, action-spaces and reward functions.
@@ D.2 Hold-out Tasks
D.3 Fine-Tuning
In Figure 15, we present the fine-tuning evaluation performance on the held-out tasks. We compare xLSTMs trained from scratched against xLSTMs initialized with the pre-trained weights. We do observe consistent improvement of the pre-trained models over the models trained from scratch. While we train on a substantial number of environments, the total amount of data used is still only a fraction of that employed in training other large-scale models, such as LLMs. Consequently, we do not observe comparable few-shot generalization. However, we anticipate that few-shot generalization capabilities will emerge as we increase both data volume and model size.
@@ D.3 Fine-Tuning
D.4 In-context Learning
We assess the ICL abilities of modern recurrent architectures on the Dark-Room environment considered in prior works on in-context RL (Laskin et al., 2022; Lee et al., 2023; Schmied et al., 2024b). In Dark-Room, the agent is located in a dark room. The task is to navigate to an invisible goal location in that dark room. The state is partially observable, as the agent only observes its own x-y position on the grid ($|\mathcal{S}|=2$). The action space consists of 5 discrete actions: move up, move down, move left, move right, stay ($|\mathcal{A}|=5$). Upon reaching the goal location, the agent receives a reward of +1 for every step in the episode it resides on the goal location. Consequently, the agent first has to explore the room to find the goal. Once the goal location is found (as indicated by the positive reward), the agent can exploit this knowledge. Given a multi-episodic context, the agent should be able to exploit information contains in the previous trials (e.g., exploiting one path vs. avoiding another).
In our experiments, the Dark-Room is a $10\times 10$ grid and episodes last for 100 steps, starting in the top left corner of the grid. We adopt the same experiment setup as Schmied et al. (2024b) and leverage their datasets. We train 16M parameter agents on datasets from 80 randomly selected goal locations in the grid. The datasets contain 100K transitions per task and are obtained by training task-specific PPO (Schulman et al., 2018) agents. Then, we evaluate the in-context abilities of our agents on 20 hold-out goal locations. During evaluation, the agent is given 40 episodes to interact with the environment, which we refer to as ICL-trials. Furthermore, we adopt the AD (Laskin et al., 2022) framework for training our agents with a multi-episodic context. We use the same sequence representation as used in our main experiments, consisting of states, returns-to-go (target return set to 80 during evaluation), and rewards. Note that this differs from the sequence representation used by Laskin et al. (2022). We set the context length for all agents to the equivalent of two episodes, which amounts to 200 timesteps in total.
In Figure 16, we report the ICL performance over the 40 ICL trials on (a) 80 training locations and (b) 20 hold-out locations for the 4 different backbones considered in this work. We observe that the recurrent backbones attain considerably higher scores than the Transformer backbone. Furthermore, we find that xLSTM [7:1] attains the highest overall scores, which we attribute to the state-tracking abilities (Merrill et al., 2024) of sLSTM blocks. We aim to explore the ICL abilities of modern recurrent backbones more in future work.
@@ D.4 In-context Learning
D.5 Inference Time Comparisons
We empirically examine the difference in inference speed between of our models. Similar to De et al. (2024), we report both latency and throughput. For real-time applications, latency is the more important dimension, and therefore we focus our analysis on latency.
D.5.1 Latency
In Figures 17 and 18, we report the latencies for DT and xLSTM with the same number of layer blocks as DT, and twice the number of layers blocks as DT, respectively. We conduct our comparison for two different batch sizes and across varying sequence lengths.
(a) $B=1$
(b) $B=16$
@@ D.5.1 Latency
D.5.2 Throughput
In Figures 19 and 20, we similarly report the attained throughput for DT and xLSTM with the same number of layer blocks as DT, and twice the number of layers blocks as DT, respectively. We conduct our comparison for two fixed context lengths and varying batch sizes.
(a) $C=800$
(b) $C=1600$
@@ D.5.2 Throughput
D.5.3 xLSTM Kernel Comparisons
@@
Appendix E Ablations
E.1 Removing action condition
E.1.1 DT on Meta-World
We found that removing actions from the context results in better performance across backbones. In Figure 22, we report the learning curves over 200K updates for DT with varying context lengths on Meta-World, both with and without actions in the context. While context lengths beyond 1 hurt performance when training with actions, the reverse is true when training without actions. This is in contrast to recent works, which did not benefit from longer contexts (Octo Model Team et al., 2024). However, while removing actions improves performance on Meta-World, it does not affect performance on discrete control. On Meta-World, we observed that the models become overly confident (high action logits), which is problematic if poor initial actions are produced. We assume this is because in robotics actions change smoothly and by observing previous actions the agent learns shortcuts. A similar issue has been identified by Wen et al. (2020), and termed the copycat problem, because the agent is incentivized to copy previous actions. Our solution is to remove actions from the input sequence. This prevents the agent from learning shortcuts and alleviates the copycat problem.
@@ E.1.1 DT on Meta-World
E.1.2 DT on all 432 tasks.
To further investigate the effect of removing actions from the context, we repeat this ablation on the full 432 tasks and 6 domains at the 206M model scale. In Figure 23, we report the learning curves for a DT with varying sequence lengths trained (a) with and (b) without actions in the agent’s context. Similar to the single-domain study on Meta-World with smaller models, we find that providing a longer context does not improve performance, resulting in a normalized score of around 0.3 across domains. In contrast, without action in the context, we observe a consistent improvement in the evaluation performance as the sequence length increases. In fact, the normalized score increase from around 0.3 with $C=1$ to 0.7 with $C=50$. For computational reasons, we only report one seed per sequence length in this experiment, but we believe that the overall trends are clear.
(a) w/ actions
(b) w/o actions
Figure 23: Ablation on removing the action condition for varying context lengths $C$. Performance of DT (a) with, and (b) without action condition on all 432 tasks. Without actions in the context, the performance of DT improves with increasing $C$.
To better understand on which domains the longer context benefits or hurts our agents, we also present the normalized score per domain in Figure 24. Without actions in the context, we find that longer context consistently benefits the performance across domains. With actions in the context we observe that on Meta-World and DMControl, the performance deteriorates for $C>1$. In contrast, on the discrete control domains Atari and Procgen, but also on the continuous continous control domain Composuite, performance tends to improve with $C>1$. This suggests that the copycat problem is particularly present on Meta-World and DMControl. However, note that the final performances on Atari, Procgen and Mimicgen are considerably worse when actions are present in the context compared to when they are not.
(a) w/ actions
(b) w/o actions
Figure 24: Ablation on removing the action condition for varying context lengths $C$. We show the normalized score per domain for all context lengths (a) with and (b) without actions.
To further investigate this, we compute the MSE between subsequent actions in the training dataset (similar to Wen et al. (2020)) for the continuous control domains and report them in Table 7. Indeed, we find that Meta-World and DMControl exhibit significantly lower MSEs between subsequent actions than Composuite. While Mimicgen also exhibits a low MSE between consecutive actions, all backbones perform poorly on this challenging benchmark. Consequently, we conclude that removing actions from the agent’s context is particularly effective for domains where actions change smoothly.
Table 7: Average MSE ($\pm$ standard deviation) between subsequent actions in robotics datasets.
Meta-World DMControl Composuite Mimicgen
Avg. MSE $0.08_{\pm 0.09}$ $0.2_{\pm 0.22}$ $2.1_{\pm 0.3}$ $0.015_{\pm 0.007}$
This result highlights the fact that large action models can strongly benefit from increased context length, even on the simulated environments we consider in this work. Furthermore, we believe that this effect can be even bigger in complex real-world environments that require longer-term interactions.
E.1.3 xLSTM on all 432 tasks.
To validate that modern recurrent backbones also benefit from training with longer sequence lengths, we repeat the same ablation as presented in Appendix E.1.2 using xLSTM [1:0]. We report the learning curves validation perplexities and evaluation performance across all 432 tasks for varying context lengths in Figure 25. Note that the validation perplexity curves in Figure 25a, start at step 50K for readability. Again, we observe considerable improvements in the validation perplexities and in the normalized scores (0.4 for $C=1$ to 0.8 for $C=50$) as the context length increases.
(a) Sequence Prediction Performance
(b) Evaluation Performance
Figure 25: Ablation on the effect of for varying the context length $C$ for xLSTM. We report (a) validation perplexity and (b) evaluation performance across the 432 training tasks for xLSTM [1:0]. Without actions in the context, the performance of DT improves with increasing $C$.
In addition, we provide the normalized scores per domain for xLSTM with varying sequence lengths in Figure 26. Across domains, we observe increasing performance with increasing $C$.
(a) w/o actions
Figure 26: Ablation on the effect of for varying the context length $C$ for xLSTM. We show the normalized scores per domain for all context lengths.
E.2 Return-conditioning vs. Behavior Cloning
Across experiments presented in the main text, except for the ICL experiments, we utilized a sequence representation that includes return-to-go tokens (RTG) as commonly used in the DT literature (Chen et al., 2021; Lee et al., 2022). At inference time, the RTG allows to condition the model on a high target return to produce high-quality actions. This is particularly useful when the datasets contain a mixture of optimal and suboptimal trajectories. However, many recent works focus on behavior cloning without return conditioning (Brohan et al., 2023b, a; Octo Model Team et al., 2024).
To better understand whether our findings transfer to the behavior cloning setting, we conduct an ablation study in which we exclude the RTG tokens and the reward tokens from the sequence representation. This means that the sequence consists of state and reward tokens, or state-tokens only. In Figures 27 and 27, we report the (a) validation perplexities and (b) evaluation performance on the 432 task for the four considered backbones when removing RTG or RTG and reward, respectively. We retain the same training settings and datasets as reported in Appendix C (200K updates, evaluation after every 50K steps). We observe similar learning dynamics as for the 206M models that include RTG/reward tokens in the sequence representation (see Figure 2 and Figure 11). Consequently, we conclude that the same performance trends holds for training the considered backbones with and without RTG/reward condition. Note, that the final performances are lower compared to the models that include the RTG condition and that can be conditioned on a high return at inference time.
(a) Sequence Prediction Performance
(b) Evaluation Performance
Figure 27: Ablation on the effect of omitting the RTG condition. We report the learning curves for (a) validation perplexity and (b) evaluation performance across the 432 training tasks for 206M parameter models. We observe similar performance trends as when including the RTG in the sequence.
(a) Sequence Prediction Performance
(b) Evaluation Performance
Figure 28: Ablation on the effect of omitting the RTG condition and the reward condition. We report the learning curves for (a) validation perplexity and (b) evaluation performance across the 432 training tasks for 206M parameter models. We observe similar performance trends as when including the RTG in the sequence.
E.3 Effect of mLSTM-to-sLSTM ratio.
Throughout our experiments, we compare two xLSTM variants: xLSTM [7:1] and xLSTM [1:0]. The bracket notation was introduced by (Beck et al., 2024), and denotes the ratio of mLSTM to sLSTM blocks. For example, xLSTM [7:1] contains 1 sLSTM block for every 7 mLSTM blocks. As described in Appendix C, we aim to maintain the same ration as proposed by Beck et al. (2024). While mLSTM blocks are fully parallelizable, sLSTM blocks are not. However, sLSTM preserves the non-diagonalized recurrent matrix to enable state-tracking (Merrill et al., 2024). As such, sLSTM can be attractive for tasks that require state-tracking (see Figure 4 in Beck et al. (2024)).
We first conduct an ablation study on the effect of the mLSTM-to-sLSTM ratio on the evaluation performance across all 432 tasks. For this experiment, we use the 16M parameter model that contains 8 xLSTM blocks in total. Consequently, we compare the following ratios [1:0] (only mLSTM), [0:1] (only sLSTM), [1:1], [1:3], [7:1]. In addition, we investigate the placement of sLSTMs across all 8 blocks. To indicate the placement, we use @ followed by the layer index (starting at 0). For example, [3:1] @ 1,3 indicates that the second and fourth layer are sLSTMs. In Figure 29 we report the validation perplexities and evaluation performance for different ratios and layer placements across the 432 tasks. For computational reasons, we conduct this experiment with only 1 seed per ratio. We find that at the 16M parameter scale, xLSTM [1:0] on average outperforms the variants that leverage sLSTM blocks. This indicates that these domains do not strongly benefit from the state tracking abilities of sLSTM.
(a) Sequence Prediction Performance
(b) Evaluation Performance
Figure 29: Ablation on the effect of the mLSTM-to-sLSTM ratio. We report the learning curves for (a) validation perplexity and (b) evaluation performance across the 432 training tasks for 206M parameter models with varying ratios.
Next, conduct the same analysis on Dark-Room $10\times 10$ ICL environment as used in Appendix D.4. Unlike most of the 432 tasks used in our main experiments, Dark-Room exhibits a partially-observable observation space and sparse rewards. Consequently, Dark-Room is more likely to require state tracking abilities. In fact, we already observed better performance for xLSTM [7:1] than for xLSTM [1:0] in Appendix 16. In Figure 30, we report the ICL curves for the 80 train tasks and 20 hold-out tasks. We observe that xLSTM variants that contain sLSTM blocks at lower-level positions, such as [7:1] @ 1 and [3:1] @ 1,3 outperform xLSTM [1:0]. In contrast, xLSTM variants that contain sLSTM blocks at deeper-level positions, such as [0:1] and 3:1 @ 5,7, perform poorly. This is similar to findings by Beck et al. (2024) who also place sLSTM layers at lower-level positions.
(a) 80 training tasks
(b) 20 hold-out tasks
Figure 30: In-context Learning on Dark-Room $10\times 10$ for varying mLSTM-to-sLSTM ratios.
We conclude that sLSTM layers can be important building blocks for tasks that require state-tracking, such as Dark-Room. Most of the 432 tasks we consider in the main experiments of this work contain fully observable observation spaces and may not require state-tracking. However, we believe that more complex tasks with longer horizons or partial observability, as is common in real-world applications, could greatly benefit from the state-tracking abilities provided by sLSTM blocks. As such, equipping an agent with the ability to perform state-tracking by including sLSTM blocks may be a valuable option for practitioners. This is a distinguishing factor of xLSTM from Mamba, which does not exhibit state-tracking.
E.4 Effect of Dropout in DT
@@
DTs use by default a Dropout (Srivastava et al., 2014) rate of 0.1. However, during our experiments, we found that Dropout has detrimental effects on the evaluation performance, particularly on continuous control domains like Composuite. In Figure 31, we show the validation perplexities and evaluation performance for a DT trained with and without Dropout. Consequently, we remove Dropout from our DT variant.
@@
E.5 Effect of reducing number of layers in xLSTM
In prior works, xLSTM and Mamba use twice the number of layers blocks as the Transformer baseline, while maintaining the same hidden dimension (Gu & Dao, 2023; Beck et al., 2024). For our inference-time comparisons, we therefore reduce the number of layer blocks in xLSTM by half. To ensure a fair comparison, we consequently adjust the hidden size of xLSTM to match the number of parameters of the Transformer baseline. In this section, we investigate the effect of these modifications of the xLSTM architecture on the model performance.
@@ E.5 Effect of reducing number of layers in xLSTM
Appendix F Embedding Space Analysis